# A Nonperturbative Toolkit for Quantum Gravity

### v2

@@ 3.1.1 Shell wormholes
The shell states are useful because the associated wormholes and their action can be explicitly computed, at least in the large shell mass limit. The wormholes are constructed by identifying shell worldvolumes across two disks subject to the Israel junction conditions as indicated in Figs. 6 and 7. The saddlepoint geometry on either side of the shell can be anything that locally solves the equations of motion; hence by symmetry these are portions of any of the saddlepoints contributing to the Z-partition function in (4). Generalizing this construction, we construct wormhole saddlepoints contributing to the amplitude $\overline{(G^{n})_{ii}}=\overline{\braket{i}{j_{1}}\braket{j_{1}}{j_{2}}\cdots\braket{j_{n}}{i}}$ (no sum on $i,j_{k}$) by connecting $n$ asymptotic boundaries (arising from each factor in $(G^{n})_{ii}$) with a single multiboundary wormhole. The $k$-th shell is able to propagate through the bulk of the wormhole from $\mathcal{O}_{k}$ to $\mathcal{O}^{\dagger}_{k}$ on different asymptotic boundaries, and hence the index structure of $\overline{(G^{n})_{ii}}$ manifests in the bulk as the wormhole in Fig. 8. We will call this wormhole geometry the pinwheel. There are also disconnected saddlepoints connecting different subsets of the boundaries together. These look like collections of smaller pinwheels.
@@
To construct a pinwheel saddlepoint imagine that we cut Fig. 8 open along the shell worldvolumes producing the so-called $L$- and $R$- sheet diagrams in Fig. 9(a). We can produce these sheets by removing the pink shell homology regions from the $L$ and $R$ disks in Fig. 9(b). The $L$ ($R$) disk boundary contains $n$ shell operator insertions $\mathcal{O}_{1,2,\cdots n},\mathcal{O}^{\dagger}_{1,2,\cdots n}$ with $\mathcal{O}_{i}$ and $\mathcal{O}^{\dagger}_{i}$ separated by a time $\Delta T_{L,i}$ ($\Delta T_{R,i}$) and $\mathcal{O}^{\dagger}_{i},\mathcal{O}_{i+1}$ separated by a time $\beta_{L}$ ($\beta_{R}$). Constructively, then, we start with the disks in Fig. 9(b), remove the pink shell homology regions, and then use the Israel junction conditions Israel:1966rt to glue along the indicated shell worldvolumes. The propagation times $\Delta T_{i}$ are determined dynamically by the junction conditions to obtain an on-shell wormhole geometry. This procedure can be applied to any saddle of the Z partition function defined above, and hence there are in general multiple saddles.
@@
As shown in Balasubramanian:2022gmo ; Balasubramanian:2022lnw ; Antonini:2023hdh the bulk turning points of the shells approach the asymptotic boundary of the disk in the large shell mass ($m_{i}\to\infty$) limit and hence the purple shell homology regions pinch off. In this limit, $\Delta T_{i}\to 0$, and each shell contributes simply contributes a factor $Z_{m_{i}}\sim e^{-2(d-1)log(\mathrm{G_{N}}m_{i})}$ that is a universal function of its mass. Furthermore, this behavior is independent of the disk saddle geometries that are being glued along the shells Balasubramanian:2022gmo ; Balasubramanian:2022lnw ; Antonini:2023hdh . Hence, in the saddle approximation the fully connected contribution to the n-boundary path integral decomposes into two Z-partition functions and a shell contribution: $\overline{(G^{n})_{ii}}|_{connected}=\overline{Z}(n\beta_{L})\overline{Z}(n\beta_{R})\prod_{i=1}^{n}Z_{m_{i}}$.
In the $m_{i}\to\infty$ limit the diagram in Fig. 6 computing the norm $\braket{i}{i}$ also simplifies – the propagation time shell propagator shrinks to zero, so that the diagram decomposes into two disks of circumference $\beta_{L,R}$ attached via a pointlike shell trajectory so that Eq. 14 for the norm becomes $\braket{i}{j}=\delta_{ij}Z_{1}=\delta_{ij}\overline{Z}(\beta_{L})\overline{Z}(\beta_{R})Z_{m_{i}}$ within the saddlepoint approximation. We can also see this by setting $n=1$ in the above expression for $\overline{(G^{n})_{ii}}|_{connected}$. If we then consider the normalized set of states |$\ket{k}/\sqrt{Z_{1}}$, we get
@@ 3.1.2 Shell geometry
Focusing now on the asymptotically AdS case for concreteness, we can associate a geometry to a shell state by considering the leading gravitational saddlepoint contributing to the norm $\overline{\langle i|i\rangle}=\overline{Z}(\beta_{L})\overline{Z}(\beta_{R})Z_{m_{i}}$ (Fig. 6). As shown above, in the large shell mass limit this is determined by the leading saddlepoint for the Z partition functions $\overline{Z}(\beta_{L}),\overline{Z}(\beta_{R})$. The leading Z-saddle depends on whether the inverse effective temperature, i.e. Euclidean time period $\beta_{L,R}$, is above or below the Hawking-Page transition Hawking:1982dh at $\beta_{HP}$. For example if $\beta_{L}>\beta_{HP}$ the $L$ disk saddle is thermal AdS, whereas for $\beta_{L}<\beta_{HP}$ it is the Euclidean AdS black hole (similarly for $R$). The time reflection symmetric bulk slice of these saddles can be analytic continued to Lorentzian signature to associate a Lorentzian geometry to the state. As $\beta_{L},\beta_{R}$ can be tuned independently there are 4 regimes to consider:
@@
    $\beta_{L},\beta_{R}<\beta_{HP}$ : The leading saddles are two-sided Schwarschild-AdS black holes with a long wormhole interior (Fig. 10(a)). We will refer to these as type 1 shell states and the space spanned by these states as $\mathcal{H}_{1}\subseteq\mathcal{H}_{Shell}$. These are the states originally considered in Balasubramanian:2022gmo .
@@
    $\beta_{L},\beta_{R}>\beta_{HP}$: The leading saddles are two copies of thermal AdS with the addition of a compact Big-Crunch AdS cosmology (Fig. 10(b)). This construction was studied in detail in Antonini:2023hdh . We will refer to these as type 2 shell states and the space spanned by them as $\mathcal{H}_{2}\subseteq\mathcal{H}_{Shell}$.
@@
    $\beta_{L}<\beta_{HP}$ and $\beta_{R}>\beta_{HP}$ or vice versa : The leading saddles correspond to an $L$ ($R$) single-sided AdS Black hole and a disconnected $R$ ($L$) copy of thermal AdS (Fig. 10(c)). We refer to these as type 3 shell states and the space spanned by them states as $\mathcal{H}_{3}\subseteq\mathcal{H}_{Shell}$. Such states were studied in Balasubramanian:2022lnw , but they did not include the disconnected thermal AdS.
@@ 3.2 Inserting a complete set of shell states
We will now argue that the shell states span $\mathcal{H}_{LR}$ at least in a coarse-grained sense. There is already some evidence for this from the argument of Balasubramanian:2022gmo ; Balasubramanian:2022lnw ; Climent:2024trz that any sufficiently large collection of shell states, projected into a microcanonical band where the dominant saddlepoint is an eternal black hole, spans a space of dimension $e^{\mathbf{S}(E_{L})+\mathbf{S}(E_{R})}$ where where $\mathbf{S}(E)$ is the Beckenstein-Hawking entropy of a horizon associated to energy $E$. However, this approach to completeness of the shell states requires an independent argument that the dimension of the full Hilbert space in this regime is $e^{\mathbf{S}(E_{L})+\mathbf{S}(E_{R})}$. There are also technical complexities involving the microcanonical projection. To develop the tools for the full argument in Sec. 4.2 and Sec. 5, we start with an alternative argument that the shell Hilbert space $\mathcal{H}_{shell}$ at fixed preparation temperature spans $\mathcal{H}_{LR}$ by showing that the projector onto $\mathcal{H}_{shell}$ (9) acts as the identity on $\mathcal{H}_{LR}$:
@@
The gravitational path integral for $\overline{Z(\beta)}$, which has a boundary manifold of topology $\mathbb{S}_{\beta}\times\mathcal{X}^{d-1}$, has a standard interpretation following Gibbons and Hawking as the thermal partition function of quantum gravity. If this is indeed the case, the state $|\beta\rangle\in\mathcal{H}_{LR}$ obtained by cutting this path integral in half (recall that $Z(\beta)=\langle\beta|\beta\rangle$) must have support on all states in $\mathcal{H}_{LR}$. Indeed, in the AdS/CFT setting the holographic dictionary equates $|\beta\rangle$ to the thermofield-double (TFD) state defined on two copies of the CFT: $|TFD(\beta)\rangle=\sum_{i}e^{\frac{-\beta E_{n}}{2}}|n\rangle_{L}|n\rangle_{R}$, which has support on all of the energy eigenstates of the theory. We can test the validity of (17) by checking whether the following equality is true:
@@ 3.2.1 Resolvent Resummation
For example, consider $Tr{G^{2}}=\sum_{i,j}\langle i|j\rangle\langle j|i\rangle=\sum_{i,j}Z_{2}+Z_{1}^{2}\delta_{ij}=\kappa^{2}Z_{2}+\kappa Z_{1}^{2}$. As the disconnected contribution requires $j=i$ it must break the $j$ index loop and is suppressed by a factor $1/\kappa$, whereas higher topology contributions, like $Z_{2}$, are suppressed by factors of $e^{-1/G_{N}}$. We are going to consider a “planar” limit in which $\kappa,e^{1/G_{N}}$ are both large while $\kappa\,e^{-1/G_{N}}\sim\mathcal{O}(1)$, discussed in Penington:2019kki ; Balasubramanian:2022gmo . In this limit both planar partially-connected and planar fully-connected geometries contribute at leading order in $1/\kappa$ and $e^{-1/\mathrm{G_{N}}}$. By planar we mean a pattern of connectivity between the asymptotic boundaries that involves no crossings when depicted on a 2D plane (see for example Sec. 5.3 in Boruch:2024kvv ). Non-planar wormhole geometries require breaking at least two index loops, and for every such geometry there is a planar geometry with the same pattern of connectivity between boundaries. The non-planar geometries are thus subleading and we will exclude them from the sum over saddles in the planar limit.
In the planar limit the sum over saddles can be performed by considering the fully-connected saddle geometries connecting two of the tadpoles to $n$ shell boundaries (Fig. 12(a)), which we refer to as a folded wormhole. We denote the sum of all folded wormhole saddles as $\tilde{Z}_{n+2}$. These folded wormholes are distinct from the pinwheels explained above and considered in Balasubramanian:2022gmo ; Balasubramanian:2022lnw ; Climent:2024trz ; Balasubramanian:2024yxk . Rather, these wormholes are generalizations of the $n=0$ case discussed in Sec. 3 of Sasieta:2022ksu .
@@
In the planar limit we can use a resolvent resummation method Penington:2019kki ; Balasubramanian:2022gmo ; Balasubramanian:2024yxk ; Boruch:2024kvv to express the sum over all planar saddles for $\overline{G^{n}_{ij}\langle\beta|i\rangle\langle j|\beta\rangle}$ in terms of just the fully-connected wormhole saddles via a Swinger-Dyson equation. In particular, we define the resolvent matrix
@@
See Boruch:2024kvv for details on the $c_{0}$ integration contour. As we discussed above, the shell boundaries have two shells each that must be matched up by grouping together the tadpole boundaries. Diagrammatically, we can write the tadpoles as in Fig. 13, leading to the schematic resummation in Fig. 14.
@@ 3.2.2 Microcanonical Projection
The sum (22) cannot be performed analytically due to the $\overline{Z}((n+1)\beta_{L}+(n+1)\beta_{R}+\beta)$ factor. Even within the leading saddle approximation, at large enough $n$ the asymptotic disk length will be above $\beta_{HP}$, causing the dominant disk saddle to shift from the large black hole to thermal AdS. We can evade these difficulties by considering instead the insertion of the projector onto the microcanonical shell Hilbert space $\mathcal{H}_{shell,E}$ at fixed $L,R$ energies $E_{L}=E_{R}=E$. As discussed in Balasubramanian:2024yxk , microcanonical shell states are obtained by Laplace transform of the canonical shell states with respect to $e^{E\beta_{L}}$ and $e^{E\beta_{R}}$. Hence from (22) we have
@@
where $\overline{Z}(\beta_{L})_{E_{L}=E}$ is defined as the Laplace transform of $\overline{Z}(\beta_{L})$ etc. The Laplace transforms of the factors in the denominator yield $\overline{Z}(\beta_{L})_{E_{L}=E}=e^{\mathbf{S}(E_{L})}$ (and similarly for $R$), whereas the numerator becomes $\overline{Z}((n+1)(\beta_{L}+\beta_{R})+\beta)_{E_{L}=E_{R}=E}=e^{-\beta E}e^{\mathbf{S}(E)}$. Explicit calculation shows that for energies above a certain threshold the black hole saddle dominates so that $\mathbf{S}(E)$ is the Bekenstein-Hawking entropy at energy $E$ with small corrections from the subleading saddles. Below this threshold $\mathbf{S}(E)=1$ because the thermal saddle dominates. Plugging this into (23) yields:
@@
The analytic structure of the shell state resolvent was discussed in detail in Balasubramanian:2024yxk . The upshot is that once $\kappa\geq e^{\mathbf{S}(E)}$ the resolvent develops a pole at $\lambda=0$ and we can deform the contour $c_{0}\rightarrow c_{2}$ which encircles $\lambda=0$ clockwise, obtaining:
@@
We have found that either the family of would-be TFD states has support only on shell states, or in the microcanonical ensemble any set of $\kappa\geq e^{\mathbf{S}(E)}$ shell states resolves the identity, i.e., (17) is true, and hence spans the complete microcanonical Hilbert space. The latter also demonstrates that the dimension of this Hilbert space is $e^{\mathbf{S}(E)}$, within the leading saddlepoint approximation, consistently with the Bekenstein-Hawking entropy formula.

### v3

@@ 3.1.1 Shell wormholes
The shell states are useful because the associated wormholes and their action can be explicitly computed, at least in the large shell mass limit. The wormholes are constructed by identifying shell worldvolumes across two disks subject to the Israel junction conditions as indicated in Figs. 6 and 7. The saddlepoint geometry on either side of the shell can be anything that locally solves the equations of motion; hence by symmetry these are portions of any of the saddlepoints contributing to the Z-partition function in (4). Generalizing this construction, we construct wormhole saddlepoints contributing to the amplitude $\overline{(G^{n})_{ii}}=\overline{\braket{i|j_{1}}\braket{j_{1}|j_{2}}\cdots\braket{j_{n}|i}}$ (no sum on $i,j_{k}$) by connecting $n$ asymptotic boundaries (arising from each factor in $(G^{n})_{ii}$) with a single multiboundary wormhole. The $k$-th shell is able to propagate through the bulk of the wormhole from $\mathcal{O}_{k}$ to $\mathcal{O}^{\dagger}_{k}$ on different asymptotic boundaries, and hence the index structure of $\overline{(G^{n})_{ii}}$ manifests in the bulk as the wormhole in Fig. 8. We will call this wormhole geometry the pinwheel. There are also disconnected saddlepoints connecting different subsets of the boundaries together. These look like collections of smaller pinwheels.
@@
To construct a pinwheel saddlepoint imagine that we cut Fig. 8 open along the shell worldvolumes producing the so-called $L$- and $R$- sheet diagrams in Fig. 9(a). We can produce these sheets by removing the pink shell homology regions from the $L$ and $R$ disks in Fig. 9(b). The $L$ ($R$) disk boundary contains $n$ shell operator insertions $\mathcal{O}_{1,2,\cdots n},\mathcal{O}^{\dagger}_{1,2,\cdots n}$ with $\mathcal{O}_{i}$ and $\mathcal{O}^{\dagger}_{i}$ separated by a time $\Delta T_{L,i}$ ($\Delta T_{R,i}$) and $\mathcal{O}^{\dagger}_{i},\mathcal{O}_{i+1}$ separated by a time $\beta_{L}$ ($\beta_{R}$). Constructively, then, we start with the disks in Fig. 9(b), remove the pink shell homology regions, and then use the Israel junction conditions Israel (1966) to glue along the indicated shell worldvolumes. The propagation times $\Delta T_{i}$ are determined dynamically by the junction conditions to obtain an on-shell wormhole geometry. This procedure can be applied to any saddle of the Z partition function defined above, and hence there are in general multiple saddles.
@@
As shown in Balasubramanian et al. (2024c, b); Antonini et al. (2023) the bulk turning points of the shells approach the asymptotic boundary of the disk in the large shell mass ($m_{i}\to\infty$) limit and hence the purple shell homology regions pinch off. In this limit, $\Delta T_{i}\to 0$, and each shell contributes simply contributes a factor $Z_{m_{i}}\sim e^{-2(d-1)log(\mathrm{G_{N}}m_{i})}$ that is a universal function of its mass. Furthermore, this behavior is independent of the disk saddle geometries that are being glued along the shells Balasubramanian et al. (2024c, b); Antonini et al. (2023). Hence, in the saddle approximation the fully connected contribution to the n-boundary path integral decomposes into two Z-partition functions and a shell contribution: $\overline{(G^{n})_{ii}}|_{connected}=\overline{Z}(n\beta_{L})\overline{Z}(n\beta_{R})\prod_{i=1}^{n}Z_{m_{i}}$.
[…]
@@ 3.1.2 Shell geometry
Focusing now on the asymptotically AdS case for concreteness, we can associate a geometry to a shell state by considering the leading gravitational saddlepoint contributing to the norm $\overline{\langle i|i\rangle}=\overline{Z}(\beta_{L})\overline{Z}(\beta_{R})Z_{m_{i}}$ (Fig. 6). As shown above, in the large shell mass limit this is determined by the leading saddlepoint for the Z partition functions $\overline{Z}(\beta_{L}),\overline{Z}(\beta_{R})$. The leading Z-saddle depends on whether the inverse effective temperature, i.e. Euclidean time period $\beta_{L,R}$, is above or below the Hawking-Page transition Hawking and Page (1983) at $\beta_{HP}$. For example if $\beta_{L}>\beta_{HP}$ the $L$ disk saddle is thermal AdS, whereas for $\beta_{L}<\beta_{HP}$ it is the Euclidean AdS black hole (similarly for $R$). The time reflection symmetric bulk slice of these saddles can be analytic continued to Lorentzian signature to associate a Lorentzian geometry to the state. As $\beta_{L},\beta_{R}$ can be tuned independently there are 4 regimes to consider:
@@
    $\beta_{L},\beta_{R}<\beta_{HP}$ : The leading saddles are two-sided Schwarschild-AdS black holes with a long wormhole interior (Fig. 10(a)). We will refer to these as type 1 shell states and the space spanned by these states as $\mathcal{H}_{1}\subseteq\mathcal{H}_{Shell}$. These are the states originally considered in Balasubramanian et al. (2024c).
@@
    $\beta_{L},\beta_{R}>\beta_{HP}$: The leading saddles are two copies of thermal AdS with the addition of a compact Big-Crunch AdS cosmology (Fig. 10(b)). This construction was studied in detail in Antonini et al. (2023). We will refer to these as type 2 shell states and the space spanned by them as $\mathcal{H}_{2}\subseteq\mathcal{H}_{Shell}$.
@@
    $\beta_{L}<\beta_{HP}$ and $\beta_{R}>\beta_{HP}$ or vice versa : The leading saddles correspond to an $L$ ($R$) single-sided AdS Black hole and a disconnected $R$ ($L$) copy of thermal AdS (Fig. 10(c)). We refer to these as type 3 shell states and the space spanned by them states as $\mathcal{H}_{3}\subseteq\mathcal{H}_{Shell}$. Such states were studied in Balasubramanian et al. (2024b), but they did not include the disconnected thermal AdS.
@@ 3.2 Inserting a complete set of shell states
We will now argue that the shell states span $\mathcal{H}_{LR}$ at least in a coarse-grained sense. There is already some evidence for this from the argument of Balasubramanian et al. (2024c, b); Climent et al. (2024) that any sufficiently large collection of shell states, projected into a microcanonical band where the dominant saddlepoint is an eternal black hole, spans a space of dimension $e^{\mathbf{S}(E_{L})+\mathbf{S}(E_{R})}$ where where $\mathbf{S}(E)$ is the Beckenstein-Hawking entropy of a horizon associated to energy $E$. However, this approach to completeness of the shell states requires an independent argument that the dimension of the full Hilbert space in this regime is $e^{\mathbf{S}(E_{L})+\mathbf{S}(E_{R})}$. There are also technical complexities involving the microcanonical projection. To develop the tools for the full argument in Sec. 4.2 and Sec. 5, we start with an alternative argument that the shell Hilbert space $\mathcal{H}_{shell}$ at fixed preparation temperature spans $\mathcal{H}_{LR}$ by showing that the projector onto $\mathcal{H}_{shell}$ (9) acts as the identity on $\mathcal{H}_{LR}$:
@@
The gravitational path integral for $\overline{Z(\beta)}$, which has a boundary manifold of topology $\mathbb{S}_{\beta}\times\mathcal{X}^{d-1}$, has a standard interpretation following Gibbons and Hawking as the thermal partition function of quantum gravity. If this is indeed the case, the state $|\beta\rangle\in\mathcal{H}_{LR}$ obtained by cutting this path integral in half (recall that $Z(\beta)=\langle\beta|\beta\rangle$) must have support on all the energy eigenstates of the theory. Indeed, in the AdS/CFT setting the holographic dictionary equates $|\beta\rangle$ to the thermofield-double (TFD) state defined on two copies of the CFT: $|TFD(\beta)\rangle=\sum_{n}e^{\frac{-\beta E_{n}}{2}}|n\rangle_{L}|n\rangle_{R}$. We can test the validity of (17) by checking whether the following equality is true:
@@ 3.2.1 Resolvent Resummation
For example, consider $Tr{G^{2}}=\sum_{i,j}\langle i|j\rangle\langle j|i\rangle=\sum_{i,j}Z_{2}+Z_{1}^{2}\delta_{ij}=\kappa^{2}Z_{2}+\kappa Z_{1}^{2}$. As the disconnected contribution requires $j=i$ it must break the $j$ index loop and is suppressed by a factor $1/\kappa$, whereas higher topology contributions, like $Z_{2}$, are suppressed by factors of $e^{-1/G_{N}}$. We are going to consider a “planar” limit in which $\kappa,e^{1/G_{N}}$ are both large while $\kappa\,e^{-1/G_{N}}\sim\mathcal{O}(1)$, discussed in Penington et al. (2022); Balasubramanian et al. (2024c). In this limit both planar partially-connected and planar fully-connected geometries contribute at leading order in $1/\kappa$ and $e^{-1/\mathrm{G_{N}}}$. By planar we mean a pattern of connectivity between the asymptotic boundaries that involves no crossings when depicted on a 2D plane (see for example Sec. 5.3 in Boruch et al. (2024a)). Non-planar wormhole geometries require breaking at least two index loops, and for every such geometry there is a planar geometry with the same pattern of connectivity between boundaries. The non-planar geometries are thus subleading and we will exclude them from the sum over saddles in the planar limit.
In the planar limit the sum over saddles can be performed by considering the fully-connected saddle geometries connecting two of the tadpoles to $n$ shell boundaries (Fig. 12(a)), which we refer to as a folded wormhole. We denote the sum of all folded wormhole saddles as $\tilde{Z}_{n+2}$. These folded wormholes are distinct from the pinwheels explained above and considered in Balasubramanian et al. (2024c, b); Climent et al. (2024); Balasubramanian et al. (2025). Rather, these wormholes are generalizations of the $n=0$ case discussed in Sec. 3 of Sasieta (2023).
@@
In the planar limit we can use a resolvent resummation method Penington et al. (2022); Balasubramanian et al. (2024c, 2025); Boruch et al. (2024a) to express the sum over all planar saddles for $\overline{G^{n}_{ij}\langle\beta|i\rangle\langle j|\beta\rangle}$ in terms of just the fully-connected wormhole saddles via a Swinger-Dyson equation. In particular, we define the resolvent matrix
@@
See Boruch et al. (2024a) for details on the $c_{0}$ integration contour. As we discussed above, the shell boundaries have two shells each that must be matched up by grouping together the tadpole boundaries. Diagrammatically, we can write the tadpoles as in Fig. 13, leading to the schematic resummation in Fig. 14.
@@ 3.2.2 Microcanonical Projection
The sum (22) cannot be performed analytically due to the $\overline{Z}((n+1)\beta_{L}+(n+1)\beta_{R}+\beta)$ factor. Even within the leading saddle approximation, at large enough $n$ the asymptotic disk length will be above $\beta_{HP}$, causing the dominant disk saddle to shift from the large black hole to thermal AdS. We can evade these difficulties by considering instead the insertion of the projector onto the microcanonical shell Hilbert space $\mathcal{H}_{shell,E_{L},E_{R}}$ at fixed $L,R$ energy windows of with $\Delta E_{L,R}$ centered around $E_{L},E_{R}$. As discussed in Balasubramanian et al. (2025), microcanonical shell states are obtained by Laplace transform of the canonical shell states. [^1] This follows by noting the Hamiltonian generates boundary time evolution and assuming the factorisation (12). In particular, we can project the two-boundary shell states into $L,R$ microcanonical windows $E_{L,R}-\frac{\Delta E_{L,R}}{2},E_{L,R}+\frac{\Delta E_{L,R}}{2}]$ by the double Laplace transform:
[…]
Performing the microcanonical projection of (22) in this way we get
@@
where we have defined $\overline{Z}(\beta_{L})_{E_{L}}\equiv e^{-E_{L}\beta_{L}+\mathbf{S}(E_{L})}$ (similarly for $R$), and $\overline{Z}((n+1)(\beta_{L}+\beta_{R})+\beta)_{E_{L},E_{R}}\equiv\delta_{E_{L},E_{R}}e^{-((n+1)(\beta_{L}+\beta_{R})E}e^{-\beta E}e^{\mathbf{S}(E)}$. The $E_{L}=E_{R}=E$ constraint in the numerator arises from the fact that the thermo-field double state $|\beta\rangle$ is diagonal in the energy basis, and for ease we have taken $\Delta{E_{L}}=\Delta{E_{R}}$. Explicit calculation shows that for energies above a certain threshold the black hole saddle dominates so that $\mathbf{S}(E)$ is the Bekenstein-Hawking entropy at energy $E$ with small corrections from the subleading saddles. Below this threshold $\mathbf{S}(E)=1$ because the thermal saddle dominates. Plugging this into (24) yields:
@@
The analytic structure of the shell state resolvent was discussed in detail in Balasubramanian et al. (2025). The upshot is that once $\kappa\geq e^{\mathbf{S}(E)}$ the resolvent develops a pole at $\lambda=0$ and we can deform the contour $c_{0}\rightarrow c_{2}$ which encircles $\lambda=0$ clockwise, obtaining:
@@
Note that if $\kappa<e^{\mathbf{S}(E)}$, evaluating $\overline{\langle\beta|\prod_{\mathcal{H}{\text{shell}}}|\beta\rangle}$ requires a detailed analysis of the functional form of the resolvent $R(\lambda)$. Only when $\kappa\geq e^{\mathbf{S}(E)}$ can the projected overlap be computed straightforwardly, yielding equation (28). We emphasize that the insertion of $G^{n}_{ij}|i\rangle\langle j|$ and the subsequent analytic continuation $n\to-1$ are therefore nontrivial operations, in contrast to a simpler insertion such as $e^{-\frac{1}{a}\hat{\mathcal{O}}}$ followed by the limit $a\to\infty$.
We have found that either the family of would-be TFD states has support only on shell states, or in the microcanonical ensemble any set of $\kappa\geq e^{\mathbf{S}(E)}$ shell states resolves the identity, i.e., (17) is true, and hence spans the complete microcanonical Hilbert space. The latter also demonstrates that the dimension of this Hilbert space is $e^{\mathbf{S}(E)}$, within the leading saddlepoint approximation, consistently with the Bekenstein-Hawking entropy formula. This result implies that the state counting performed in Balasubramanian et al. (2024b, c); Climent et al. (2024); Balasubramanian et al. (2024a) indeed measures the dimension of the full (semiclassical) microcanonical Hilbert space, rather than that of some restricted “shell superselection sector”.